# Towards Graph Foundation Models: Learning Generalities Across Graphs via Task-Trees

### v2

@@ F.2 Domain: E-Commerce
Node Classification. The comprehensive node classification results on e-commerce datasets are presented in Table 15. Our proposed GIT model outperforms the baselines in most settings, particularly for the specialized version. Specialization significantly improves performance in zero-shot and in-context learning, highlighting the advantages of using task-trees as the basic learning instances. In the basic fine-tuning setting, we also observe that supervised methods (GCN and GAT) generally outperform self-supervised methods, such as GraphMAE (Hou et al., 2022) and OFA (Liu et al., 2024), indicating the occurrence of negative transfer. However, GIT surpasses these supervised methods on 5 out of 7 datasets, further demonstrating the benefits of task-trees as basic learning instances. It it important to note that we consider Children and Ratings as heterophily graphs (Chen et al., 2024b) due to their low homophily ratio.

### v3

@@ F.2 Domain: E-Commerce
Node Classification. The comprehensive node classification results on e-commerce datasets are presented in Table 17. Our proposed GIT model outperforms the baselines in most settings, particularly for the specialized version. Specialization significantly improves performance in zero-shot and in-context learning, highlighting the advantages of using task-trees as the basic learning instances. In the basic fine-tuning setting, we also observe that supervised methods (GCN and GAT) generally outperform self-supervised methods, such as GraphMAE (Hou et al., 2022) and OFA (Liu et al., 2024a), indicating the occurrence of negative transfer. However, GIT surpasses these supervised methods on 5 out of 7 datasets, further demonstrating the benefits of task-trees as basic learning instances. It it important to note that we consider Children and Ratings as heterophily graphs (Chen et al., 2024b) due to their low homophily ratio.